OpenStudy (anonymous):

Solve the system of equations. (Enter your answers from smallest to largest value of the imaginary part of x. Enter NONE in any unused answer blanks.) x + y = 16 xy = 68

OpenStudy (anonymous):

First, in one of the equations, isolate either x or y.

OpenStudy (anonymous):

Then, plug in the isolated variable into the other equation, solve for the other variable.

OpenStudy (anonymous):

i tried this and it isnt looking right.. i dont know what i am doing wrong

OpenStudy (anonymous):

Can you tell me what you've done so far?

OpenStudy (anonymous):

x + y = 16 xy = 68 Let's work with equation 1: x = 16 - y Now that we have one variable isolated, use this in the other equation. xy = 68 (16 - y)y = 68 16y - y^2 = 68 Solve for y. After you solve for y, plug y back into your isolated x: x = 16 - y Now you should have x and y.

OpenStudy (anonymous):

ok let me work on it.

OpenStudy (anonymous):

i know its a simple problem but i keep getting the wrong answer

OpenStudy (anonymous):

Do you have answer choices?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

so can you help me?

OpenStudy (anonymous):

What was your answer?

OpenStudy (anonymous):

for y= 8+- 2i

OpenStudy (anonymous):

then I inserted this in x... and I dont know if I am solving it wrong... but I got... -44

OpenStudy (anonymous):

y = 8 - 2i is correct for x you should get x = 8 + 2i

OpenStudy (anonymous):

ok so I was multiplying.. i dont even know how I got that... ok so, for x2... do u get that when u do the negative form of y?

OpenStudy (anonymous):

yes, i just saw that

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

thank you!

OpenStudy (shinalcantara):

you forgot to divide 4i by 2

OpenStudy (shinalcantara):

yw =)

OpenStudy (anonymous):

that paper was my first round, i went at it three rounds.. when i get tired, i dont process math well

OpenStudy (shinalcantara):

i wrote it down. my typing skills is really bad and it would take me years to finish typing lol.. =]

OpenStudy (anonymous):

i understood perfectly, thank u